OpenStudy (anonymous):

3. A 12 kg barrel is pulled up by a rope. The barrel accelerates at 1.2 m/s2. Find the force exerted by the rope. Show all your work.

OpenStudy (anonymous):

there is only y direction movement yes? so let's ignore sum F forces in the x direction = 0 since that's a given sum F forces in the y direction: What are the forces acting on the barrel? the rope (T) and the weight (mg) of the barrel. sum forces in the y ---> T -mg = ma we want rope force: T = ma + mg plug chug! T = m(a-g) = 12(1.2+9.8) I do not have a calculation on me but the units of the number is in Newtons(N)

OpenStudy (anonymous):